OpenStudy (anonymous):

1+sinx+sin^2x+sin^3x+.... = 4+2√3 Solve for x

OpenStudy (anonymous):

\[1+sinx+\sin^2x+\sin^3x+.... = 4+2√3\]

OpenStudy (anonymous):

0<x<pi, x is not equal to pi/2 then find x

hartnn (hartnn):

do you find any pattern in that infinite series ? hint : geometric sequence

OpenStudy (anonymous):

OMG! :P That just didn't strike me! Shoot! Lemme try that!

OpenStudy (anonymous):

hey @hartnn , just one thing.. shall i take 1 as the first term of sinx .. I mean, 1/sinx = r if 1 =a and r=sinx if a=sinx

hartnn (hartnn):

a is the first term of the sequence. here, first term = 1 so a = 1

hartnn (hartnn):

r= sin x

hartnn (hartnn):

so you should get \(\large \dfrac{1}{1-\sin x} = 4+2\sqrt3\) solve this for sin x